1. Plot the graph of the equations 3x - 8y = 5 and 4x - 2y = 11 and interpret the result.
Answer by midwood_trail(310)

I will provide the steps for what to do and you can do the work.
STEPS:
1-Graph each equation on the xy-plane.
2-Find the point where the two graphs meet or cross.
You can easily see the point of intersection when using graph paper.
Go get some graph paper. You can find free graphing paper online.
If the two graphs meet, there will be one and ONLY ONE result or one solution.
If the two graphs do not meet, there is no solution for this system of equations in two variables.
Understand?
I received your reply about not understanding the steps provided.
Plot the graph of the equations 3x - 8y = 5 and 4x - 2y = 11 and interpret the result.
I will show you how to graph one of the equations and you can follow my steps and solve the other equation. How does that sound?

3x - 8y = 5
1-we isolate y.
-8y = -3x + 5
2-divide both sides by -8 to isolate y.
y = (-3x + 5)/-8
y = (3x/8) - (5/8)....This can be written as y = (3x - 5)/8

3-select 4 values (or more values) for x to find y values and make your points.

Let x = -1
y = (3(-1)- 5)/8
y = -8/8
y = -1
First point to plot: (-1, -1).
=====================
Let x = 0
y = (3(0) - 5)/8
y = -5/8
Our second point to plot: (0, -5/8)
=====================
Let x = 1
y = (3(1) - 5)/8
y = -2/8
y = -1/4
Our third point to plot: (1,-1/4)
======================
Let x = 2
y = (3(2) - 5)/8
y = 1/8
Our fourth point to plot: (2, 1/8)
Graph these points on the xy-plane.
Then do the same thing with the other equation given.
After plotting about 4 points (or more) for the other equation, look for a piece on both graphs where they meet or cross.
This "meeting place" on the xy-plane will be your solution.
Got it?